OpenStudy (anonymous):

In 1990, there were 10.8 thousand radio stations on the air in the United States. In 1998, there were 12.6 thousand. What was the average rate of change per year in the number of radio stations in the United States from 1990 to 1998?

OpenStudy (anonymous):

The total number of radio stations added divided by the total number of years will give you the average rate of change per year.

OpenStudy (anonymous):

12.6 - 10.8 = 1.8 thousand radio stations added 1998 - 1990 = 10 years 1800/10 = 180 radio stations per year

OpenStudy (anonymous):

satellite could i use you for one question?